OpenStudy (alicia.marie018):

A scientist counts 35 bacteria present in a culture and finds the the number of bacteria triples each hour. The function y= 35×3^x models the number of bacteria after x hours. A. Graph the function B. Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (aaronandyson):

This one will be: Given function \[y = 35 \times 3^x\] When the count of bacteria is 550 in the culture the function will be: \[550 = 35 \times 3^x\] Now divide 550 by 35 \[\frac{ 550 }{ 35 } = 3^x\] Divide the L.H.S(Or the Left-hand side). by 5. 550/5 = 110 35/5 = 7 \[\frac{ 550 }{ 35 } = 3^x\] when divided by 5 gives \[\frac{ 110 }{ 7 } = 3^x\]

OpenStudy (aaronandyson):

\[\frac{ 550 }{ 35 } = 3^x\] equals \[\frac{ 110 }{ 7 } = 3^x\] equals \[x = \ln (\frac{ 110 }{ 7 } ) /\ln (3)\] then use a calculator.

OpenStudy (aaronandyson):

ln(110 / 7) / ln(3) = 2.50731786 In approximately 2.51 hours.

OpenStudy (alicia.marie018):

Thank you

OpenStudy (aaronandyson):

Where is my medal..?